# Theory of Valley Splitting in Si/SiGe Spin-Qubits: Interplay of Strain, Resonances and Random Alloy Disorder

### v1

@@ II.5.2 Random Component
where the integrals $J_{n}^{\mathrm{rand}}$ are given in Eq. (66) and the band structure coefficients $C_{n}^{\left(4\right)}$ are defined in Eq. (67). We refer to Appendix B for details. The band structure coefficients $C_{n}^{\left(4\right)}$ are plotted as a function of shear strain in Fig. 2 (b). In contrast to the deterministic part, non-trivial resonances are less pronounced in the disorder-induced contribution to the valley splitting such that Eq. (35) is in fact largely dominated by the trivial contribution proportional to $C_{n=0}^{\left(4\right)}$. In general, modifications due to shear strain are weaker for the random component than for the deterministic part.
@@ II.5.3 Valley Splitting Statistics
From the normal distribution of $\Delta_{\mathrm{rand}}$ it immediately follows that the valley splitting obeys a Rice distribution
@@
where $I_{n}\left(x\right)$ denotes the modified Bessel functions of first kind [62]. The asymptotics
@@
Tab. 1: Parameter values used in the computations, if not stated otherwise. Band structure parameters were computed as a function of strain using the empirical pseudopotential model described in Ref. [49] with parameters stated there.
@@ III Results
Fig. 3: Impact of interface width on valley splitting. (a) Expected valley splitting $\langle E_{\mathrm{VS}}\rangle$ and shape parameters $\nu$ and $\sigma$ of the Rice distribution in the absence of shear strain $\varepsilon_{x,y}=0$ as a function of the QW interface width, cf. Eq. (10). The interface width is given in units of monolayers (ML) $a_{0}/4\approx 0.1357\,\mathrm{nm}$ of relaxed Si. The gray shaded region indicates the deterministically enhanced regime according to Eq. (39) and the red shaded region shows the [25%, 75%] quantile of the Rice distribution. Deterministic enhancements are observed for sharp interfaces with up to one ML width. (b) Same as (a), but for small shear strain $\varepsilon_{x,y}=0.1\%$. The deterministically enhanced regime is extended to about four MLs. (c) Absolute values of the individual components $\Delta_{\mathrm{det},n}$ contributing to the deterministic valley splitting, see Eq. (27), as a function of the interface width. In the absence of shear strain, the resulting valley slitting is dominated by the $n=0$ and $n=-2$ resonances in the deterministically enhanced regime, whereas the $n=0$ resonance dominates in the disorder-dominated regime. (d) In the case of small shear strain $\varepsilon_{x,y}=0.1\%$ the $n=-1$ component is the dominant contribution over the entire parameter domain. In the plot, an electric field of $F=10\,\mathrm{mV}/\mathrm{nm}$ was assumed.
@@ III.2 Unconventional Heterostructures
where $X_{\mathrm{QW}}=X_{b}\left(1-\Xi\left(z\right)\right)$ is the profile of the conventional SiGe/Si/SiGe quantum well, see Eq. (9), and
@@ III.2.1 Uniform Germanium Concentration
For $q=0$, the profile in Eq. (40), reduces to a uniform Ge concentration in the QW. In this configuration, which was first proposed in Ref. [17], the mean valley splitting is enhanced because the random Ge concentration corresponds to a flat (white noise) power spectrum in reciprocal space, contributing even at large wave numbers. The increase of the valley splitting is, however entirely disorder-dominated. Therefore, numerous spin-valley hotspots must be expected [26].
For a sufficient amount of shear strain, when the deterministic component starts to be dominated by the $n=-1$ resonance, cf. Fig. 3 (d), slight enhancements of the deterministic contribution can be expected. These enhancements are, however, much smaller than the disorder-induced component for typical parameters. This is shown in Fig. 4 (a)–(b), where the valley splittings at $q=0$ are practically unchanged even in the presence of shear strain. Moreover, Fig. 5 (b) shows a regime with enhanced mean valley splitting at very low wave numbers, which is clearly outside of the deterministically enhanced regime (indicated by the separatrix condition (39), dashed line). Finally, the typical epitaxial profile and electron density distribution is shown in Fig. 5 (c) along with the power spectral density (PSD) of the product $\left(U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)\right)\psi^{2}_{0}\left(z\right)$ and the complex-plane distribution of the intervalley coupling parameter $\Delta$, which further evidences the disorder-dominated nature of this concept.
@@ III.2.2 Short-Period Wiggle-Well
The results shown in Fig. 4 feature a strong resonance at $q=2k_{0}$, which grows with increasing Ge amplitude $X_{\mathrm{ww}}$, but is practically independent of shear strain. This configuration is called the short-period wiggle-well. The effect is based on a direct enhancement of the coupling between the valley states within the same Brillouin zone triggered by the periodicity of the confinement potential. The short-period wiggle-well requires a high-frequent Ge modulation with a very short wave length $\lambda=\pi/k_{0}\approx 2.4\,\mathrm{ML}\approx 0.32\,\mathrm{nm}$. The epitaxial growth of such rapidly modulated structures is out of reach with the current technology, which is why the short-period wiggle-well not relevant for practical applications.
@@
Fig. 5: (a) Mean valley splitting of the long-period wiggle-well with fixed wave number $q=0.32\times 2\pi/a_{0}\approx 2k_{1}$ as a function of shear strain $\varepsilon_{x,y}$ and Ge amplitude $X_{\mathrm{ww}}$. The dashed line is the separatrix defined in Eq. (39) that separates disorder-dominated and deterministically enhanced regimes. For large shear strains, a deterministic enhancement is achieved already for very small Ge amplitudes. The red line indicates the constant shear strain value considered in panel (b). (b) Mean valley splitting as a function of wave number $q$ and amplitude $X_{\mathrm{ww}}$ for fixed shear strain $\varepsilon_{x,y}=0.07\%$. Three domains with valley splitting enhancements can be observed, namely, at very low $q\approx 0$ (neary constant uniform Ge concentration in the QW), near $q\approx k_{1}$ (lower harmonic/ Ge spike, only for sufficiently high $X_{\mathrm{ww}}$) and near $q\approx 2k_{1}$ (long-period wiggle-well). The dashed line is again the separatrix. (c) Left: Epitaxial profile and ground state electron density distribution for the uniform Ge concentration in the well ($q=0$, $X_{\mathrm{ww}}=0.05$) at shear strain $\varepsilon_{x,y}=0.07\%$. The ratio $\nu/\left(2\sigma\right)$ indicates that this configuration falls into the disorder-dominated regime. Right: Power spectral density (PSD) of the product $\left(U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)\right)\psi_{0}^{2}\left(z\right)$ as a function of the wave number $k$. The contribution to the valley splitting is dominated by the $2k_{1}$-resonance and is thus larger than predicted by the $2k_{0}$-theory. The inset shows the statistical distribution of the intervalley coupling parameter in the complex plane in the range of $-0.5\,\mathrm{meV}\leq\mathrm{Re}\left(\Delta\right),\mathrm{Im}\left(\Delta\right)\leq 0.5\,\mathrm{meV}$. (d) Same as (c), but at ($q=k_{1}$, $X_{\mathrm{ww}}=0.17$). The product of the potential and the wave function yields a strong Fourier component at $k=2k_{1}$, despite the potential has only half of the frequency $q=k_{1}$. The configuration has a moderate deterministic enhancement $\nu/\left(2\sigma\right)\approx 0.55$. (e) Same as (c), but for a long-period wiggle-well at ($q=2k_{1}$, $X_{\mathrm{ww}}=0.05$). The $q=2k_{1}$-periodic modulation of the potential provides a significant deterministic enhancement of the valley splitting $\nu/\left(2\sigma\right)\approx 0.86$. A constant electric field $F=5\,\mathrm{mV}/\mathrm{nm}$ and biaxial (tensile) strain was assumed in all simulations.
@@ III.2.3 Long-Period Wiggle-Well
In the presence of shear strain, a second resonance emerges at $q=2k_{1}$, see Fig. 4 (b), where $k_{1}$ has been defined in Eq. (30). The peak grows linearly with increasing amplitude $X_{\mathrm{ww}}$. This resonance is attributed to the so-called long-period wiggle-well [33, 30, 32], which enhances the coupling of valley states in neighboring Brillouin zones separated by $2k_{1}$. The corresponding wave length is $\lambda\approx 12.4\,\mathrm{ML}\approx 1.68\,\mathrm{nm}$. Epitaxial growth of such structures is within reach with molecular beam epitaxy.
@@
The two-parametric plot in Fig. 5 (a) shows the dependency of the mean valley splitting in a long-period wiggle-well on shear strain and the Ge amplitude. The separatrix (dashed line) indicates that a minimum shear strain of about $\varepsilon_{x,y}\approx 0.02\%$ is required to enter the deterministically enhanced regime even for high Ge amplitudes. For large shear strains above $\varepsilon_{x,y}\gtrsim 0.08\%$, however, the deterministic enhancement sets in already for tiny Ge amplitudes (but then also with very small mean valley splittings). In Fig. 5 (b), the mean valley splitting $\langle E_{\mathrm{VS}}\rangle$ is shown for fixed shear strain $\varepsilon_{x,y}=0.07\%$ as a function of the wave number $q$ and the Ge oscillation amplitude $X_{\mathrm{ww}}$. The broad resonance around $q\approx 2k_{1}$ falls clearly in the deterministically enhanced regime and is relatively robust against slight deviations from the ideal wave number. The typical epitaxial profile and the corresponding electron density distribution are shown in Fig. 5 (e) along with the PSD and the complex-plane distribution of the intervalley coupling parameter (inset). The strong spectral contribution at $2k_{1}$ leads to a strong deterministic enhancement of the valley splitting such that the complex normal distribution governing the statistical properties is steered away from the origin. Hence, a significant suppression of spin-valley hotspots can be expected. The long-period wiggle-well is therefore a very promising approach to achieve large, deterministic enhancements of the valley splitting.
@@ III.2.4 Lower Harmonic Wiggle-Well / Ge-Spike
While at first glance the longer wave length might be favorable due to reduced demands on the epitaxial growth process, the required high Ge content causes severe drawbacks. In fact, we expect a strong enhancement of the spin-orbit interaction in such a structure [64], which would greatly complicate the control of the spin qubit. In addition, the deterministic enhancements achieved with such a structure are lower than that of a corresponding long-period wiggle-well.
Finally, we remark that the second Ge-peak near the lower interface is not essential for the observed effect. It is simply enforced by the parametrization in Eq. (40). Without the second Ge-peak, the structure has strong similarities with the Ge-spike described in Ref. [34], but with a much broader Ge-peak.
@@ IV Discussion
The model presented in this paper differs in a number of aspects from the model by Feng & Joynt [30], which has been the first to provide an explanation for the mechanism behind the long-period wiggle-well in the framework of envelope function theory. The mechanism proposed there to induce a violation of the nonsymmorphic screw symmetry – which otherwise suppresses the resonance at $2k_{1}$ – is based on a randomization of the plane wave expansion coefficients of the Bloch factors. The latter are determined for statistical ensembles of disordered SiGe alloys using a combination of an empirical pseudopotential model and an extended virtual crystal approximation, whereby the alloy disorder induces the desired symmetry breaking. This approach is in principle very plausible, but a subsequent analysis shows that the corresponding statistical distribution of the valley splitting implied by this method does not lead to a Rice distribution (as observed in experiments and atomistic simulations). Instead, a significantly more skewed distribution is found, which has a significantly larger probability density at low valley splittings compared to the Rice distribution. Furthermore, in the model by Feng & Joynt, the distribution of the intervalley coupling parameter in the complex plane takes a strongly elliptical shape, which is very different from the nearly circular distribution obtained here, see Appendix C.1. Another consequence of the symmetry-breaking mechanism proposed there is that the variance of the valley splitting in the long-period wiggle-well becomes very large. Thus, the valley splitting enhancement due to the long-period wiggle-well is found to be primarily disorder-dominated and is by no means deterministic. This is reasonable when one assumes that the symmetry breaking is disorder-induced. These predictions on the characteristics of the long-period wiggle-well are in sharp contrast to those obtained in the present work, where the nonsymmorphic screw symmetry is broken via shear strain. As a consequence, the present model predicts a strong deterministic enhancement of the valley splitting in the long-period wiggle-well. The actual microscopic symmetry-breaking mechanism, however, is very likely a combination of both effects. The impact of disorder on the long-period wiggle-well might therefore require further investigation. Finally, the model by Feng & Joynt shows another pronounced resonance for a wiggle well with intermediate wave number $q=k_{0}$. In our model, we found that this is caused by a harmonic similar to the configuration described in Sec. III.2.4. For typical SiGe parameters (recall that Feng & Joynt used a conduction band offset more typical for Si-MOS), however, the small deterministic resonance is entirely obscured by the disorder-induced component. Therefore, no peak can be observed in the mean valley splitting in Fig. 4.
@@
Finally, we discuss the relation between the present work and the recent paper by Woods et al. [32]. In Ref. [32], the shear strain induced symmetry-breaking mechanism to unlock the resonance behind the long-period wiggle-well was studied for the first time. The envelope function model presented there, which is derived from a one-dimensional tight-binding model, features an intervalley coupling matrix element with unconventional functional form that is notably different from other expressions in the literature. In the present model described in Sec. II of this work, however, the standard form of the intervalley matrix element is preserved even in the presence of strain effects. The numerical results obtained here are qualitatively in an excellent agreement with those reported by Woods et al.. In particular the joint linear dependence of the long-period wiggle-well on shear strain and Ge concentration is recovered using the pseudopotential model. Furthermore, we also find good quantitative agreements in the magnitude of the obtained valley splittings. The quantitative comparison requires caution, however, as the present model also contains disorder-induced contributions, which are omitted in Ref. [32].

### v2

@@ II.5.2 Random Component
where the integrals $J_{n}^{\mathrm{rand}}$ are given in Eq. (66) and the band structure coefficients $C_{n}^{\left(4\right)}$ are defined in Eq. (67). We refer to Appendix B for details. The band structure coefficients $C_{n}^{\left(4\right)}$ are shown as a function of shear strain in Fig. 2 (b). In contrast to the deterministic part, non-trivial resonances are less pronounced in the disorder-induced contribution to the valley splitting such that Eq. (35) is in fact largely dominated by the trivial contribution proportional to $C_{n=0}^{\left(4\right)}$. In general, modifications due to shear strain are weaker for the random component than for the deterministic part.
@@ II.5.3 Valley Splitting Statistics
From the normal distribution of $\Delta_{\mathrm{rand}}$ it follows that the valley splitting obeys a Rice distribution
@@
where $I_{n}\left(x\right)$ denotes the modified Bessel functions of first kind. The asymptotics
@@
Tab. 1: Parameter values used in the computations, if not stated otherwise. Band structure parameters (including effective masses) were computed as a function of strain using the empirical pseudopotential model described in Ref. [51] with parameters stated there. The QW thickness is given in units of monolayers $\mathrm{ML}=a_{0}/4$ of the relaxed Si crystal.
@@ III Results
Fig. 3: Impact of interface width on valley splitting. (a) Mean valley splitting $\langle E_{\mathrm{VS}}\rangle$ and shape parameters $\nu$ and $\sigma$ of the Rice distribution in the absence of shear strain $\varepsilon_{x,y}=0$ as a function of the QW interface width, cf. Eq. (10). The interface width is given in units of monolayers $\mathrm{ML}=a_{0}/4\approx 0.1357\,\mathrm{nm}$ of relaxed Si. The gray shaded region indicates the deterministically enhanced regime according to Eq. (39) and the red shaded region shows the [25%, 75%] quantile of the Rice distribution. Deterministic enhancements are observed for sharp interfaces with up to one ML width. (b) Same as (a), but for small shear strain $\varepsilon_{x,y}=0.1\%$. The deterministically enhanced regime is extended to about four MLs. (c) Absolute values of the individual components $\Delta_{\mathrm{det},n}$ contributing to the deterministic valley splitting, see Eq. (27), as a function of the interface width. In the absence of shear strain, the resulting valley slitting is dominated by the $n=0$ and $n=-2$ resonances in the deterministically enhanced regime, whereas the $n=0$ resonance dominates in the disorder-dominated regime. (d) In the case of small shear strain $\varepsilon_{x,y}=0.1\%$ the $n=-1$ component is the dominant contribution over the entire parameter domain. In the simulation, an electric field of $F=10\,\mathrm{mV}/\mathrm{nm}$ was assumed.
@@ III.2 Unconventional Heterostructures
where $X_{\mathrm{QW}}=X_{b}\left(1-\Xi\left(z\right)\right)$ is the profile of the conventional SiGe/Si/SiGe QW, see Eq. (9), and
@@ III.2.1 Uniform Germanium Concentration
For $q=0$, the profile in Eq. (40), reduces to a uniform Ge concentration in the QW. In this configuration, which was first proposed in Ref. [18], the mean valley splitting is enhanced because the random Ge concentration corresponds to a flat (white noise) power spectrum in reciprocal space, contributing even at large wave numbers. The corresponding of the valley splitting enhancement is, however, entirely disorder-dominated. Therefore, numerous spin-valley hotspots must be expected [27].
For a sufficient amount of shear strain, when the deterministic component starts to be dominated by the $n=-1$ resonance, cf. Fig. 3 (d), slight enhancements of the deterministic contribution can be expected. These enhancements are, however, much smaller than the disorder-induced component for typical parameters. This is shown in Fig. 4 (a)–(b), where the valley splittings at $q=0$ are practically unchanged even in the presence of shear strain. Moreover, Fig. 5 (b) shows a regime with enhanced mean valley splitting at very low wave numbers, which is clearly outside of the deterministically enhanced regime (indicated by the separatrix condition (39), dashed line). Finally, the typical epitaxial profile and electron density distribution is shown in Fig. 5 (c) along with the power spectral density (PSD) of the product $\left(U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)\right)\psi^{2}_{0}\left(z\right)$ and the complex-plane distribution of the intervalley coupling parameter $\Delta$, showing that the valley splitting is indeed disorder-dominated.
@@ III.2.2 Short-Period Wiggle-Well
The results shown in Fig. 4 feature a strong resonance at $q=2k_{0}$, which grows with increasing Ge amplitude $X_{\mathrm{ww}}$, but is practically independent of shear strain. This configuration is called the short-period wiggle-well. The effect is based on a direct enhancement of the coupling between the valley states within the same Brillouin zone triggered by the periodicity of the confinement potential. The short-period wiggle-well requires a high-frequency Ge modulation with a very short wave length $\lambda=\pi/k_{0}\approx 2.4\,\mathrm{ML}\approx 0.32\,\mathrm{nm}$. The epitaxial growth of such rapidly modulated structures is out of reach with current technology, which is why the short-period wiggle-well is not relevant for practical applications.
@@
Fig. 5: (a) Mean valley splitting of the long-period wiggle-well with fixed wave number $q=0.32\times 2\pi/a_{0}\approx 2k_{1}$ as a function of shear strain $\varepsilon_{x,y}$ and Ge amplitude $X_{\mathrm{ww}}$. The dashed line is the separatrix defined in Eq. (39) that separates disorder-dominated and deterministically enhanced regimes. For large shear strains, a deterministic enhancement is achieved already for very small Ge amplitudes. The red line indicates the constant shear strain value considered in panel (b). (b) Mean valley splitting as a function of wave number $q$ and amplitude $X_{\mathrm{ww}}$ for fixed shear strain $\varepsilon_{x,y}=0.07\%$. Three domains with valley splitting enhancements can be observed, namely, at very low $q\approx 0$ (neary constant uniform Ge concentration in the QW), near $q\approx k_{1}$ (lower harmonic/ Ge spike, only for sufficiently high $X_{\mathrm{ww}}$) and near $q\approx 2k_{1}$ (long-period wiggle-well). The dashed line is again the separatrix. (c) Left: Epitaxial profile and ground state electron density distribution for the uniform Ge concentration in the well ($q=0$, $X_{\mathrm{ww}}=0.05$) at shear strain $\varepsilon_{x,y}=0.07\%$. The ratio $\nu/\left(2\sigma\right)$ indicates that this configuration falls into the disorder-dominated regime. Right: Power spectral density (PSD) of the product $\left(U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)\right)\psi_{0}^{2}\left(z\right)$ as a function of the wave number $k$. The contribution to the valley splitting is dominated by the $2k_{1}$-resonance and is thus larger than predicted by the $2k_{0}$-theory. The inset shows the statistical distribution of the intervalley coupling parameter in the complex plane in the range of $-0.5\,\mathrm{meV}\leq\mathrm{Re}\left(\Delta\right),\mathrm{Im}\left(\Delta\right)\leq 0.5\,\mathrm{meV}$. (d) Same as (c), but at ($q=k_{1}$, $X_{\mathrm{ww}}=0.17$). The product of the potential and the wave function yields a strong Fourier component at $k=2k_{1}$, despite the potential having only half of the frequency $q=k_{1}$. The configuration has a moderate deterministic enhancement $\nu/\left(2\sigma\right)\approx 0.55$. (e) Same as (c), but for a long-period wiggle-well at ($q=2k_{1}$, $X_{\mathrm{ww}}=0.05$). The $q=2k_{1}$-periodic modulation of the potential provides a significant deterministic enhancement of the valley splitting $\nu/\left(2\sigma\right)\approx 0.86$. A constant electric field $F=5\,\mathrm{mV}/\mathrm{nm}$ and biaxial (tensile) strain was assumed in all simulations.
@@ III.2.3 Long-Period Wiggle-Well
In the presence of shear strain, a second resonance emerges at $q=2k_{1}$, see Fig. 4 (b), where $k_{1}$ has been defined in Eq. (30). The peak grows linearly with increasing amplitude $X_{\mathrm{ww}}$. This resonance is associated with the so-called long-period wiggle-well [34, 31, 33], which enhances the coupling of valley states in neighboring Brillouin zones separated by $2k_{1}$. The corresponding wave length is $\lambda\approx 12.4\,\mathrm{ML}\approx 1.68\,\mathrm{nm}$. Epitaxial growth of such structures is feasible using molecular beam epitaxy.
@@
The two-parametric plot in Fig. 5 (a) shows the dependency of the mean valley splitting in a long-period wiggle-well on shear strain and the Ge amplitude. The separatrix (dashed line) indicates that a minimum shear strain of about $\varepsilon_{x,y}\approx 0.02\%$ is required to enter the deterministically enhanced regime even for high Ge amplitudes. For larger shear strains above $\varepsilon_{x,y}\gtrsim 0.08\%$, however, the deterministic enhancement sets in already for tiny Ge amplitudes (but then also with very small mean valley splittings). In Fig. 5 (b), the mean valley splitting $\langle E_{\mathrm{VS}}\rangle$ is shown for fixed shear strain $\varepsilon_{x,y}=0.07\%$ as a function of the wave number $q$ and the Ge oscillation amplitude $X_{\mathrm{ww}}$. The broad resonance around $q\approx 2k_{1}$ falls clearly in the deterministically enhanced regime and is relatively robust against deviations from the ideal wave number. The typical epitaxial profile and the corresponding electron density distribution are shown in Fig. 5 (e) along with the PSD and the complex-plane distribution of the intervalley coupling parameter (inset). The strong spectral contribution at $2k_{1}$ leads to a strong deterministic enhancement of the valley splitting such that the complex normal distribution governing the statistical properties is steered away from the origin. Hence, a significant suppression of spin-valley hotspots can be expected. The long-period wiggle-well is therefore a very promising approach to design qubits with deterministically large valley splittings.
@@ III.2.4 Lower Harmonic Wiggle-Well / Ge-Spike
While at first glance the longer wave length might be favorable due to reduced demands on the epitaxial growth process, the required high Ge content causes severe drawbacks. In fact, we expect a strong enhancement of the spin-orbit interaction in such a structure [65], which would significantly complicate the control of the qubit. The deterministic enhancements achieved with such a structure are lower than that of a corresponding long-period wiggle-well.
Finally, we remark that the second Ge-peak near the lower interface is not essential for the observed effect. It is simply enforced by the parametrization in Eq. (40). Without the second Ge-peak, the structure has strong similarities with the Ge-spike described in Ref. [35].
@@ IV Discussion
The model presented in this paper differs in a number of aspects from the model by Feng & Joynt [31], which has been the first to provide an explanation for the mechanism behind the long-period wiggle-well in the framework of envelope function theory. The mechanism proposed there to induce a violation of the nonsymmorphic screw symmetry – which otherwise suppresses the resonance at $2k_{1}$ – is based on a randomization of the plane wave expansion coefficients of the Bloch factors. The latter are determined for statistical ensembles of disordered SiGe alloys using a combination of an empirical pseudopotential model and an extended virtual crystal approximation, whereby the alloy disorder induces the desired symmetry breaking. This approach is in principle very plausible, but a subsequent analysis shows that the corresponding statistical distribution of the valley splitting implied by this method does not lead to a Rice distribution (as observed in experiments and atomistic simulations). Instead, a significantly more skewed distribution is found, which has a significantly larger probability density at low valley splittings compared to the Rice distribution. Furthermore, in the model by Feng & Joynt, the distribution of the intervalley coupling parameter in the complex plane takes a strongly elliptical shape, which is very different from the nearly circular distribution obtained here (see Appendix C.1). Another consequence of the symmetry-breaking mechanism proposed there is that the variance of the valley splitting in the long-period wiggle-well becomes very large. Thus, the valley splitting enhancement due to the long-period wiggle-well is found to be primarily disorder-dominated and is by no means deterministic. This is reasonable when one assumes that the symmetry breaking is disorder-induced. These predictions on the characteristics of the long-period wiggle-well are in sharp contrast to those obtained in the present work, where the nonsymmorphic screw symmetry is broken via shear strain. As a consequence, the present model predicts a strong deterministic enhancement of the valley splitting in the long-period wiggle-well. The actual microscopic symmetry-breaking mechanism, however, is very likely a combination of both effects. The impact of disorder on the long-period wiggle-well might therefore require further investigation. Finally, the model by Feng & Joynt shows another pronounced resonance for a wiggle well with intermediate wave number $q=k_{0}$. In our model, we found that this is caused by a harmonic similar to the configuration described in Sec. III.2.4. For typical SiGe parameters (recall that Feng & Joynt used a conduction band offset more typical for Si-MOS), however, the small deterministic resonance is entirely obscured by the disorder-induced component. Therefore, no peak can be observed in the mean valley splitting in Fig. 4.
@@
Finally, we discuss the relation between the present work and the recent paper by Woods et al. [33]. In Ref. [33], the shear strain induced symmetry-breaking mechanism to unlock the resonance behind the long-period wiggle-well has been studied for the first time. The envelope function model presented there, which is derived from a one-dimensional tight-binding model, features an intervalley coupling matrix element with unconventional functional form that is notably different from other expressions in the literature. In the present model described in Sec. II of this work, however, the standard form of the intervalley matrix element is preserved even in the presence of strain effects. The numerical results obtained here are qualitatively in an excellent agreement with those reported by Woods et al.. In particular the joint linear dependence of the long-period wiggle-well on shear strain and Ge concentration is recovered using the empirical pseudopotential model. Furthermore, we also find good quantitative agreements in the magnitude of the obtained valley splittings. The quantitative comparison, however, requires caution, as the present model also contains disorder-induced contributions, which are omitted in Ref. [33].